A bottle of wine with a cork cost $2.10. A bott of wine without a cork costs $2.00 more than the cork. What is the price of the cork?
Answer by ankor@dixie-net.com(12706)

A bottle of wine with a cork cost $2.10. A bottle of wine without a cork costs $2.00 more than the cork. What is the price of the cork?
:
Let x = cost of wine
Let y = cost of the cork
:
Write a simple equation for each phrase
:
"A bottle of wine with a cork cost $2.10."
x + y = 2.10
:
"A bottle of wine without a cork costs $2.00 more than the cork"
x = y + 2
:
"What is the price of the cork?"
:
Substitute (y+2) for x in the first equation:
x + y = 2.10
(y+2) + y = 2.10
2y + 2 = 2.10
2y = 2.10 - 2
y = .10/2
y = .05
:
Check solution:
x + y = 2.10
x + .05 = 2.10
x = 2.10 - 05
x = 2.05 which is $2 more the the cork